# Eyal Michaeli

### v1

@@ 1 Introduction
Recent generative methods evaluated for FGVC augmentation typically use real images as guidance in an Img2Img manner [16, 54, 21, 60]. While this helps maintain visual similarity to the target domain, it limits the degree of diversity that can be introduced, resulting in a trade-off between class fidelity and diversity [17] (see Figure˜1). We aim to free the generative process from this constraint of adhering to specific source images. To this end, we propose SaSPA: Structure and Subject Preserving Augmentation, a method that conditions the generation on more abstract representations rather than direct image inputs. Specifically, we leverage structural conditioning in the diffusion model via edge maps extracted from source images. This allows the generated samples to respect the broad shape and composition of objects in the target domain. Crucially, the lack of specific image conditioning enables greater flexibility in rendering surface details. To further ensure the preservation of fine-grained class characteristics, we integrate subject representation conditioning. By combining edge-based structural conditioning with category-level conditioning, SaSPA can generate highly diverse, class-consistent synthetic images without being overly influenced by any specific real data sample.
@@ 3 Method
Our goal is to augment a labeled training dataset for FGVC to increase its diversity while faithfully representing the sub-classes. The key insight of our method is to minimize reliance on any particular source image during generation and instead condition the generation on more abstract representations, thereby increasing diversity while accurately representing the designated class (see Figure˜1). To achieve this, we employ abstract conditions such as edges, which capture the object’s structure, and subject representation, which aims to preserve fine-grained class characteristics. The process, illustrated in Figure˜2, unfolds in five steps, outlined below:
@@ 3.1 Construction of Prompts
Prompt Generation via GPT-4 [1]. We input the meta-class to GPT-4 with the instruction to produce 100 unique, relevant, and diverse prompts, each inherently containing the term of the meta-class. This strategy ensures that the generated images stay true to the fundamental aspects of the meta-class while hopefully containing relevant and diverse scenarios. To increase the specificity and relevance of these prompts, we integrate the relevant sub-class into each prompt whenever it is used. Unlike a recent work [16], which also uses GPT-4 to create prompts, we do not require image captions of the dataset. The exact instructions for GPT-4 and more example prompts are in Section˜F.1.
@@ 3.2 Visual Prior Extraction
To ensure that the generated images maintain the overall structure and shape of objects belonging to their respective classes, we condition the synthesis process on edge representations extracted from real images in the dataset. Concretely, for a dataset of $N$ images, we extract one-channel edge representations using the Canny edge detector [6]. This yields a set of $N$ edge-based visual priors that capture the structural characteristics of each sub-class. By conditioning the generative model on these edge maps, we can preserve object shape and layout during synthesis while allowing flexibility in rendering other surface-level details. In Table˜15, we additionally explore the use of HED [63] as an alternative technique for edge extraction and structural conditioning.
@@ 3.3 Image Generation
Diversity in synthetic data is crucial for effective training [55, 45, 31]. Unlike most recent approaches [21, 16, 54], our method focuses on edges and subject representation as a prior rather than the source image. We show in Figure˜1 that this approach maintains class fidelity, and as a result of not using the source image, affords the model greater flexibility to introduce novel semantic features such as weather, lighting, or even new elements both within and outside the object’s confines. This strategy might be particularly beneficial in FGVC tasks, where the subtle differences between classes are crucial, and hence, maintaining class fidelity while introducing diversity is of paramount importance.
@@
Using BLIP-diffusion with ControlNet requires a prompt, an edge map, and a reference image of the target subject. To maintain sub-class accuracy while introducing diversity at the sub-class level, the reference image is selected from the same sub-class but differs from the real image used to extract the edge map (we experiment with BLIP-diffusion inputs in Table˜4). Specifically, we generate $M=2$ augmentations for each real image in the training set: we extract an edge map for each real image, and for each edge map, we randomly select $M$ prompts and $M$ subject reference images from the same sub-class. These inputs are then fed into the generation model of ControlNet with BLIP-diffusion as the base model to produce $M$ augmentations of the real image. Example augmentations are visualized at Figure˜3. DTD [9] has only one prompt per real image, because we utilize image captions as prompts for it, as explained in Appendix˜E.
@@ 3.4 Filtering
Predictive Confidence Filtering. To ensure each augmentation faithfully represents its designated sub-class, we implement a predictive confidence filtering strategy inspired by recent work [21] strategy CLIP Filtering. This method employs CLIP [41] to filter out images that do not strongly correlate with the textual labels of their class among all classes in the dataset. However, the limitation of using CLIP in this context is its insufficient granularity in understanding fine-grained concepts. For our method, we discard any augmented images for which the true label does not rank within the top-k predictions of a baseline model trained on the original dataset. This approach helps to exclude images that likely misrepresent the source sub-class, thus maintaining a high-quality dataset for model training. In our implementation, we use $k=10$. Further details about this method, the baseline model used, and other filtering techniques are discussed in Section˜F.2.
@@ 4.1 Experimental Setup
For generation, we employ BLIP-diffusion for SaSPA and Stable Diffusion v1.5 [46] for all other diffusion-based augmentation methods. For training, we follow the implementation strategy outlined in the CAL study [44], tailored for FGVC. Our models, based on the ResNet50 [20] architecture, are fine-tuned for each dataset. More data generation and training details can be found in Appendix˜E.
Comparison Methods. We benchmark our method, SaSPA, against established traditional and generative data augmentation techniques. In the traditional category, our comparisons include: CAL-Aug [44]: utilizes random flipping, cropping, and color-space variations. RandAugment [11]: applies a series of random image transformations such as rotation, shearing, and color variations to training images. CutMix [66]: generates mixed samples by randomly cutting and pasting patches between training images to encourage the model to learn more localized and discriminative features. Combined Methods: Tests the synergistic effects of CAL-Aug with CutMix and RandAug with CutMix. In the generative category, we compare with: Real-Guidance [21]: applies Img2Img with a low translation strength ($s=0.15$) to maintain high fidelity to the original images. ALIA [16]: Uses real image captions and GPT-generated domain descriptions based on these captions as prompts for Img2Img translations. Detailed descriptions of these baseline methods are in Section˜E.3.
@@ 4.2 Fine-grained Visual Classification
Datasets. We evaluate on five FGVC datasets, using the full datasets for training. We use Aircraft [30], Stanford Cars [24], CUB [58], DTD [9], and CompCars [64]. For datasets lacking a predefined validation split, we establish one. For CompCars, we utilize the exterior car parts split, focusing exclusively on classifying images of car components: head light, tail light, fog light, and front into the correct car type. Further details on the exact splits are provided in Appendix˜D.
@@
Results. We present the test accuracy of various augmentation methods in Table˜1. For each dataset, the most effective traditional augmentation method (marked by an underline) is identified using its validation set and consistently combined with all generative approaches to optimize performance for that dataset. This approach is grounded in findings that standalone generative methods generally perform better when integrated with traditional augmentations [60], a trend also evident in Table˜9.
@@ 4.3 Few-shot Learning
The training and data generation approaches remain consistent with those described in Appendix˜E, with two modifications: we use 100 epochs (down from 140), and we do not employ predictive confidence filtering for shots 4 and 8. The latter adjustment is due to the reduced reliability of the model’s predictions, resulting from the limited training data. Additionally, we increase the augmentation ratio to $\alpha=0.6$, as identified to be better for scenarios with limited data in Table˜5.
@@
Results. The results in Figure˜4 show that SaSPA consistently outperforms all other augmentation methods across all datasets and various shot counts. As seen in other works [54, 21], the benefit of augmentation diminishes as the number of shots increases, a trend most noticeable in the Cars dataset. Contrary to prior work, the gains provided by SaSPA remain substantial even at higher shot counts; notably, in the Cars dataset at 16 shots, SaSPA achieves an accuracy of 91.0%, surpassing the second-best performance of 88.3% by RG [21]. Interestingly, SaSPA sometimes matches or exceeds the performance enhancement achieved by increasing the dataset size. For example, in the DTD dataset, utilizing SaSPA with 8 shots results in an accuracy of 54.8%, slightly surpassing the 54.6% obtained when adding 50% more real data (a total of 12 shots) when relying solely on real data and the best traditional augmentation.
@@
Experimental Setup. To evaluate the effectiveness of our method in mitigating real-world contextual biases, we use the contextual bias split of the Aircraft [30] dataset constructed by Dunlap et al. [16]. The split uses two visually similar classes: Boeing-767 and Airbus-322. Each image in this split is categorized as “sky”, “grass”, or “road” depending on its background, with ambiguous examples filtered out. The bias in the dataset is introduced by training on 400 samples where Airbus aircraft are exclusively associated with road backgrounds and Boeing aircraft with grass backgrounds, although both types may appear against sky backgrounds. The exact split breakdowns are in Table˜17.
We follow the same training and generation implementation settings as for the FGVC setting (Appendix˜E), and we compare against the same generative methods. We also compare against the optimal traditional augmentation for the Aircraft dataset (CAL-Aug), as defined in Section˜4.2.
@@
Results. The results in Table˜2 show that SaSPA outperforms all other methods in overall and out-of-domain (OOD) accuracy, demonstrating its effectiveness in mitigating contextual bias. However, it falls short in in-domain (ID) accuracy. A distinct inverse relationship is observed between ID and OOD accuracy: methods that induce more significant changes from the original image—such as ALIA, which uses stronger translations than Real-Guidance (RG)—tend to achieve higher OOD accuracy but lower ID accuracy. This trend suggests that greater modifications can help reduce over-fitting to in-domain characteristics, enhancing the model’s ability to generalize effectively to new, unseen conditions. As depicted in Figure˜1, even a higher translation strength ($s=0.5/0.75$ ) yields limited diversity compared to our method. Consequently, the alterations produced by RG and ALIA are insufficient to significantly mitigate the contextual bias present in the dataset, as effective background variation is crucial for addressing such biases.
@@ 4.5 Comparing SaSPA with Concurrent Work diff-mix
Experimental Setup. In this analysis, we evaluate the performance of our SaSPA augmentation method using the diff-mix training setup, as detailed in their work. By using their open-source implementation, we further assess the robustness of our method with a different training setup. More details regarding training setup are in Section˜C.1.
@@
Results. Our results, detailed in Table˜3, highlight where SaSPA performs well and identify areas for potential improvement. The findings can be summarized as follows: (1) While diff-mix employs computationally intensive fine-tuning techniques to enhance class representation, we prioritize simplicity and lower computational demands in our approach. Despite this, SaSPA consistently outperforms diff-mix on both the Aircraft and Cars datasets across all architectures, whether combined with CutMix or used alone, demonstrating its robustness across various augmentation contexts. This also shows that conditioning the generation on more abstract representations, as we do for correct class representation, can overcome the absence of extensive fine-tuning. (2) The CUB dataset posed unique challenges, with diff-mix outperforming SaSPA using ResNet50 and both diff-mix and SaSPA under-performing relative to CutMix using ViT-B/16. Notably, despite SaSPA outperforming all methods, including CutMix in Table˜1, it does not perform as well here. We hypothesize that the use of higher resolution emphasizes finer details in each class, which may be overwhelming for SaSPA on some datasets but less so for diff-mix, likely due to the heavy fine-tuning process integrated into their method. These results indicate that some form of fine-tuning might be advantageous for complex datasets like CUB to achieve better performance. The full table, including comparisons to more augmentation methods, can be found in Section˜C.1.
@@ 5.1 Effect of Different Generation Strategies on Performance
In Table˜4, we conduct an extensive ablation study to evaluate the effectiveness of our proposed generation strategies. Specifically, we examine the integration of edge guidance, Img2Img as an alternative for edge guidance with strength = $0.5$ and in combination with edge guidance with strength = $0.85$, and subject representation. We also investigate the effect of using the same image for both edges extraction and the subject reference image (“Edges=Ref.”). Additionally, we test the impact of appending half of the prompts with artistic styles (column ‘Art.’) as described in Section˜C.5.
@@ 5.2 Effect of Augmentation Ratio on performance
In Figure˜5, we evaluate various augmentation ratios (the probability of a real image being replaced by a synthetic one in a given mini-batch). We find that for most datasets, excluding CUB, the optimal range for $\alpha$ lies between 0.2 and 0.5, with marginal differences within this range. Consequently, we selected $\alpha=0.4$ as the default augmentation ratio. However, for the CUB dataset, the default choice is $\alpha=0.1$. Considering the relatively lower improvement on CUB in Table˜1 and the underperformance on the diff-mix benchmark (Table˜3), both of which are likely attributed to lower class-fidelity, it seems that lower class fidelity necessitates a lower augmentation ratio. A possible explanation is that higher augmentation ratios are more likely to introduce bias during training when class fidelity is lower.
@@ 5.3 Effect of Augmentation Ratio on Performance with Different Amounts of Real Data
In Table˜5, we examine the interaction between $\alpha$ and the percentage of real data used. We use two $\alpha$ values for each dataset: the default value used throughout the paper and a higher value ($\alpha_{\text{high}}=\alpha+0.2$). We observe that for all amounts of real data, SaSPA achieves notable improvements with diminishing returns, similar to the trends observed in Section˜4.3. An interesting pattern emerges: as the amount of real data decreases, the optimal value of $\alpha$ tends to increase. This trend is consistent across all datasets. For instance, in the Cars dataset, when all real data is used (Frac. 1.0), $\alpha=0.6$ performs worse than $\alpha=0.4$. However, for smaller fractions of real data (e.g., 10%, 30%), using $\alpha=0.6$ yields better performance. This pattern is also observed in the Aircraft and CUB datasets, indicating that higher values of $\alpha$ are more beneficial when the amount of real data is limited.

### v2

@@ 1 Introduction
Recent generative methods evaluated for FGVC augmentation typically use real images as guidance in an Img2Img manner [16, 54, 21, 60]. While this helps maintain visual similarity to the target domain, it limits the degree of diversity that can be introduced, resulting in a trade-off between class fidelity and diversity [17] (see Figure 1). We aim to free the generative process from this constraint of adhering to specific source images. To this end, we propose SaSPA: Structure and Subject Preserving Augmentation, a method that conditions the generation on more abstract representations rather than direct image inputs. Specifically, we leverage structural conditioning in the diffusion model via edge maps extracted from source images. This allows the generated samples to respect the broad shape and composition of objects in the target domain. Crucially, the lack of specific image conditioning enables greater flexibility in rendering surface details. To further ensure the preservation of fine-grained class characteristics, we integrate subject representation conditioning. By combining edge-based structural conditioning with category-level conditioning, SaSPA can generate highly diverse, class-consistent synthetic images without being overly influenced by any specific real data sample.
@@ 3 Method
Our goal is to augment a labeled training dataset for FGVC to increase its diversity while faithfully representing the sub-classes. The key insight of our method is to minimize reliance on any particular source image during generation and instead condition the generation on more abstract representations, thereby increasing diversity while accurately representing the designated class (see Figure 1). To achieve this, we employ abstract conditions such as edges, which capture the object’s structure, and subject representation, which aims to preserve fine-grained class characteristics. The process, illustrated in Figure 2, unfolds in five steps, outlined below:
@@ 3.1 Construction of Prompts
Prompt Generation via GPT-4 [1]. We input the meta-class to GPT-4 with the instruction to produce 100 unique, relevant, and diverse prompts, each inherently containing the term of the meta-class. This strategy ensures that the generated images stay true to the fundamental aspects of the meta-class while hopefully containing relevant and diverse scenarios. To increase the specificity and relevance of these prompts, we integrate the relevant sub-class into each prompt whenever it is used. Unlike a recent work [16], which also uses GPT-4 to create prompts, we do not require image captions of the dataset. The exact instructions for GPT-4 and more example prompts are in Section F.1.
@@ 3.2 Visual Prior Extraction
To ensure that the generated images maintain the overall structure and shape of objects belonging to their respective classes, we condition the synthesis process on edge representations extracted from real images in the dataset. Concretely, for a dataset of $N$ images, we extract one-channel edge representations using the Canny edge detector [6]. This yields a set of $N$ edge-based visual priors that capture the structural characteristics of each sub-class. By conditioning the generative model on these edge maps, we can preserve object shape and layout during synthesis while allowing flexibility in rendering other surface-level details. In Table 15, we additionally explore the use of HED [63] as an alternative technique for edge extraction and structural conditioning.
@@ 3.3 Image Generation
Diversity in synthetic data is crucial for effective training [55, 45, 31]. Unlike most recent approaches [21, 16, 54], our method focuses on edges and subject representation as a prior rather than the source image. We show in Figure 1 that this approach maintains class fidelity, and as a result of not using the source image, affords the model greater flexibility to introduce novel semantic features such as weather, lighting, or even new elements both within and outside the object’s confines. This strategy might be particularly beneficial in FGVC tasks, where the subtle differences between classes are crucial, and hence, maintaining class fidelity while introducing diversity is of paramount importance.
@@
Using BLIP-diffusion with ControlNet requires a prompt, an edge map, and a reference image of the target subject. To maintain sub-class accuracy while introducing diversity at the sub-class level, the reference image is selected from the same sub-class but differs from the real image used to extract the edge map (we experiment with BLIP-diffusion inputs in Table 4). Specifically, we generate $M=2$ augmentations for each real image in the training set: we extract an edge map for each real image, and for each edge map, we randomly select $M$ prompts and $M$ subject reference images from the same sub-class. These inputs are then fed into the generation model of ControlNet with BLIP-diffusion as the base model to produce $M$ augmentations of the real image. Example augmentations are visualized at Figure 3. DTD [9] has only one prompt per real image, because we utilize image captions as prompts for it, as explained in Appendix E.
@@ 3.4 Filtering
Predictive Confidence Filtering. To ensure each augmentation faithfully represents its designated sub-class, we implement a predictive confidence filtering strategy inspired by recent work [21] strategy CLIP Filtering. This method employs CLIP [41] to filter out images that do not strongly correlate with the textual labels of their class among all classes in the dataset. However, the limitation of using CLIP in this context is its insufficient granularity in understanding fine-grained concepts. For our method, we discard any augmented images for which the true label does not rank within the top-k predictions of a baseline model trained on the original dataset. This approach helps to exclude images that likely misrepresent the source sub-class, thus maintaining a high-quality dataset for model training. In our implementation, we use $k=10$. Further details about this method, the baseline model used, and other filtering techniques are discussed in Section F.2.
@@ 4.1 Experimental Setup
For generation, we employ BLIP-diffusion for SaSPA and Stable Diffusion v1.5 [46] for all other diffusion-based augmentation methods. For training, we follow the implementation strategy outlined in the CAL study [44], tailored for FGVC. Our models, based on the ResNet50 [20] architecture, are fine-tuned for each dataset. More data generation and training details can be found in Appendix E.
Comparison Methods. We benchmark our method, SaSPA, against established traditional and generative data augmentation techniques. In the traditional category, our comparisons include: CAL-Aug [44]: utilizes random flipping, cropping, and color-space variations. RandAugment [11]: applies a series of random image transformations such as rotation, shearing, and color variations to training images. CutMix [66]: generates mixed samples by randomly cutting and pasting patches between training images to encourage the model to learn more localized and discriminative features. Combined Methods: Tests the synergistic effects of CAL-Aug with CutMix and RandAug with CutMix. In the generative category, we compare with: Real-Guidance [21]: applies Img2Img with a low translation strength ($s=0.15$) to maintain high fidelity to the original images. ALIA [16]: Uses real image captions and GPT-generated domain descriptions based on these captions as prompts for Img2Img translations. Detailed descriptions of these baseline methods are in Section E.3.
@@ 4.2 Fine-grained Visual Classification
Datasets. We evaluate on five FGVC datasets, using the full datasets for training. We use Aircraft [30], Stanford Cars [24], CUB [58], DTD [9], and CompCars [64]. For datasets lacking a predefined validation split, we establish one. For CompCars, we utilize the exterior car parts split, focusing exclusively on classifying images of car components: head light, tail light, fog light, and front into the correct car type. Further details on the exact splits are provided in Appendix D.
@@
Results. We present the test accuracy of various augmentation methods in Table 1. For each dataset, the most effective traditional augmentation method (marked by an underline) is identified using its validation set and consistently combined with all generative approaches to optimize performance for that dataset. This approach is grounded in findings that standalone generative methods generally perform better when integrated with traditional augmentations [60], a trend also evident in Table 9.
@@ 4.3 Few-shot Learning
The training and data generation approaches remain consistent with those described in Appendix E, with two modifications: we use 100 epochs (down from 140), and we do not employ predictive confidence filtering for shots 4 and 8. The latter adjustment is due to the reduced reliability of the model’s predictions, resulting from the limited training data. Additionally, we increase the augmentation ratio to $\alpha=0.6$, as identified to be better for scenarios with limited data in Table 5.
@@
Results. The results in Figure 4 show that SaSPA consistently outperforms all other augmentation methods across all datasets and various shot counts. As seen in other works [54, 21], the benefit of augmentation diminishes as the number of shots increases, a trend most noticeable in the Cars dataset. Contrary to prior work, the gains provided by SaSPA remain substantial even at higher shot counts; notably, in the Cars dataset at 16 shots, SaSPA achieves an accuracy of 91.0%, surpassing the second-best performance of 88.3% by RG [21]. Interestingly, SaSPA sometimes matches or exceeds the performance enhancement achieved by increasing the dataset size. For example, in the DTD dataset, utilizing SaSPA with 8 shots results in an accuracy of 54.8%, slightly surpassing the 54.6% obtained when adding 50% more real data (a total of 12 shots) when relying solely on real data and the best traditional augmentation.
@@
Experimental Setup. To evaluate the effectiveness of our method in mitigating real-world contextual biases, we use the contextual bias split of the Aircraft [30] dataset constructed by Dunlap et al. [16]. The split uses two visually similar classes: Boeing-767 and Airbus-322. Each image in this split is categorized as “sky”, “grass”, or “road” depending on its background, with ambiguous examples filtered out. The bias in the dataset is introduced by training on 400 samples where Airbus aircraft are exclusively associated with road backgrounds and Boeing aircraft with grass backgrounds, although both types may appear against sky backgrounds. The exact split breakdowns are in Table 17.
We follow the same training and generation implementation settings as for the FGVC setting (Appendix E), and we compare against the same generative methods. We also compare against the optimal traditional augmentation for the Aircraft dataset (CAL-Aug), as defined in Section 4.2.
@@
Results. The results in Table 2 show that SaSPA outperforms all other methods in overall and out-of-domain (OOD) accuracy, demonstrating its effectiveness in mitigating contextual bias. However, it falls short in in-domain (ID) accuracy. A distinct inverse relationship is observed between ID and OOD accuracy: methods that induce more significant changes from the original image—such as ALIA, which uses stronger translations than Real-Guidance (RG)—tend to achieve higher OOD accuracy but lower ID accuracy. This trend suggests that greater modifications can help reduce over-fitting to in-domain characteristics, enhancing the model’s ability to generalize effectively to new, unseen conditions. As depicted in Figure 1, even a higher translation strength ($s=0.5/0.75$ ) yields limited diversity compared to our method. Consequently, the alterations produced by RG and ALIA are insufficient to significantly mitigate the contextual bias present in the dataset, as effective background variation is crucial for addressing such biases.
@@ 4.5 Comparing SaSPA with Concurrent Work diff-mix
Experimental Setup. In this analysis, we evaluate the performance of our SaSPA augmentation method using the diff-mix training setup, as detailed in their work. By using their open-source implementation, we further assess the robustness of our method with a different training setup. More details regarding training setup are in Section C.1.
@@
Results. Our results, detailed in Table 3, highlight where SaSPA performs well and identify areas for potential improvement. The findings can be summarized as follows: (1) While diff-mix employs computationally intensive fine-tuning techniques to enhance class representation, we prioritize simplicity and lower computational demands in our approach. Despite this, SaSPA consistently outperforms diff-mix on both the Aircraft and Cars datasets across all architectures, whether combined with CutMix or used alone, demonstrating its robustness across various augmentation contexts. This also shows that conditioning the generation on more abstract representations, as we do for correct class representation, can overcome the absence of extensive fine-tuning. (2) The CUB dataset posed unique challenges, with diff-mix outperforming SaSPA using ResNet50 and both diff-mix and SaSPA under-performing relative to CutMix using ViT-B/16. Notably, despite SaSPA outperforming all methods, including CutMix in Table 1, it does not perform as well here. We hypothesize that the use of higher resolution emphasizes finer details in each class, which may be overwhelming for SaSPA on some datasets but less so for diff-mix, likely due to the heavy fine-tuning process integrated into their method. These results indicate that some form of fine-tuning might be advantageous for complex datasets like CUB to achieve better performance. The full table, including comparisons to more augmentation methods, can be found in Section C.1.
@@ 5.1 Effect of Different Generation Strategies on Performance
In Table 4, we conduct an extensive ablation study to evaluate the effectiveness of our proposed generation strategies. Specifically, we examine the integration of edge guidance, Img2Img as an alternative for edge guidance with strength = $0.5$ and in combination with edge guidance with strength = $0.85$, and subject representation. We also investigate the effect of using the same image for both edges extraction and the subject reference image (“Edges=Ref.”). Additionally, we test the impact of appending half of the prompts with artistic styles (column ‘Art.’) as described in Section C.5.
@@ 5.2 Effect of Augmentation Ratio on performance
In Figure 5, we evaluate various augmentation ratios (the probability of a real image being replaced by a synthetic one in a given mini-batch). We find that for most datasets, excluding CUB, the optimal range for $\alpha$ lies between 0.2 and 0.5, with marginal differences within this range. Consequently, we selected $\alpha=0.4$ as the default augmentation ratio. However, for the CUB dataset, the default choice is $\alpha=0.1$. Considering the relatively lower improvement on CUB in Table 1 and the underperformance on the diff-mix benchmark (Table 3), both of which are likely attributed to lower class-fidelity, it seems that lower class fidelity necessitates a lower augmentation ratio. A possible explanation is that higher augmentation ratios are more likely to introduce bias during training when class fidelity is lower.
@@ 5.3 Effect of Augmentation Ratio on Performance with Different Amounts of Real Data
In Table 5, we examine the interaction between $\alpha$ and the percentage of real data used. We use two $\alpha$ values for each dataset: the default value used throughout the paper and a higher value ($\alpha_{\text{high}}=\alpha+0.2$). We observe that for all amounts of real data, SaSPA achieves notable improvements with diminishing returns, similar to the trends observed in Section 4.3. An interesting pattern emerges: as the amount of real data decreases, the optimal value of $\alpha$ tends to increase. This trend is consistent across all datasets. For instance, in the Cars dataset, when all real data is used (Frac. 1.0), $\alpha=0.6$ performs worse than $\alpha=0.4$. However, for smaller fractions of real data (e.g., 10%, 30%), using $\alpha=0.6$ yields better performance. This pattern is also observed in the Aircraft and CUB datasets, indicating that higher values of $\alpha$ are more beneficial when the amount of real data is limited.